# Magic wavelength at 477 nm for the strontium clock transition

### v1

@@ Abstract
We report the experimental measurement of a magic wavelength at 476.82362(8) nm for the 88Sr clock transition. The magic wavelength is determined through AC-Stark shift spectroscopy of atoms in an optical dipole trap. The value slightly deviates from the theoretical prediction by 0.061(54) nm. This magic wavelength, being shorter than the common one at 813 nm, will be important for applications such as Bragg pulses for matter-wave interferometry involving both clock states. This work also paves the way for quantum simulation with a shorter lattice constant.
@@ I Background
Laser-cooled neutral atoms are studied for various applications, such as quantum simulation and computing Bloch et al. (2012); Henriet et al. (2020), precision sensing Geiger et al. (2020), and metrology Ludlow et al. (2015). These applications usually require optical trapping of atoms and coherent control of a transition between two internal states under consideration. Common trapping techniques include large optical dipole traps, optical lattices, and optical tweezers. With sufficiently deep optical lattices and tweezers, it is possible to suppress Doppler broadening and recoil shifts, allowing for high precision spectroscopy. However, within the trap, spatial inhomogeneity generally reduces the coherence between the states. A common technique to suppress this dephasing mechanism consists in tuning the optical lattice wavelength to the magic wavelength, canceling the differential AC-Stark shift Katori et al. (2003); Takamoto and Katori (2003); Ye et al. (2008).
@@
The magic-wavelength value depends on the atomic species and the transition under consideration. Here, ultra-narrow clock transitions of alkaline-earth atoms have attracted a lot of interest for quantum technological and precision measurement applications. Optical lattice clocks Takamoto et al. (2005); Lemke et al. (2009); Bothwell et al. (2019); Aeppli et al. (2024) that operate on these transitions are among the most precise and accurate clocks and could play an important role in a future redefinition of the second Dimarcq et al. (2024). For the clock transition of strontium atoms at 698 nm, the optical lattice generally operates at a magic wavelength of 813 nm Takamoto and Katori (2003). 813 nm magic-wavelength laser beams have also been used in strontium tweezer array systems Covey et al. (2019); Schine et al. (2022) and quantum gas microscopy Buob et al. (2024), contributing to applications of the clock transition in quantum computation Tsai et al. (2025) and quantum simulation Kolkowitz et al. (2017).
@@
For a given transition, there could be more than one magic wavelength. Finding new magic wavelengths can facilitate some specific usages. For example, the recent measurement of a magic wavelength close to 461 nm for the strontium intercombination line Kestler et al. (2022) led to progress in the state-insensitive trapping of atoms close to a nanofiber Kestler et al. (2023). For the strontium clock transition, we expect three magic wavelengths that are red detuned from the main transition at 461 nm, namely with positive real part of the polarizability. So far, only the magic wavelength at 813 nm has been used, mainly because the photon scattering rate is low, leading to potentially very long coherence times. Recently, a shorter magic wavelength at 497 nm was identified Kestler et al. (2025). Shorter wavelengths can be useful to create optical lattice with smaller inter-trap spacing. In addition, magic wavelength Bragg pulses, used in matter-wave interferometers operating on both clock states Li (2023), can benefit from shorter wavelengths that transfer larger momenta to the atoms.
@@
In this work, we experimentally measure the shortest red-detuned magic wavelength for the strontium clock transition. This magic wavelength, which is at approximately 477 nm, lies between the 474 nm transition and the 483 nm transition. As the closest red-detuned magic wavelength to the 461 nm transition, the atomic polarizability is the largest among the red-detuned magic wavelengths as shown in Fig. 1(a). The magic wavelength is determined by spectroscopic measurements of the frequency shift of the 698 nm clock transition in the presence of an optical dipole trap, and compared to a theoretical estimation.
@@
This article is organized as follows. In Sec. II, we present the theoretical estimation of the magic wavelength. In Sec. III, we discusses the experimental setup and the procedure for the spectroscopic measurements. In Sec. IV, we report the results on the 477 nm magic-wavelength measurement. In addition, we compare the differential AC-Stark-shift measurements with a model that includes the residual inhomogeneous broadening of the clock transition line. In Sec. V, we summarize our studies and provide an outlook for the uses of the 477 nm magic wavelength.
@@ II Polarizabilities estimation
The AC-stark shift of a state in the electric dipole approximation reads
@@
Generally, the polarizability can be decomposed into three contributions: the scalar, the vector, and the tensor polarizability. However, due to the spherical symmetry of the bosonic strontium clock states, the vector and tensorial terms vanish. Thus, the polarizability take the following polarization-independent form,
where is the angular frequency of the bare transition , is the electric dipole operator, and the Planck constant. The equation clearly indicates that the transitions with nearby resonant frequencies give the dominant contribution. These transitions are shown in Fig. 1(b) and their respective properties are listed in Table 1. With Eq. (2) and Table 1, we compute the polarizabilities of the clock states around 477 nm, as shown in Fig. 2.
@@ III Experiment
The experiment is performed on an ultracold cloud of 88Sr bosonic isotope. A schematic drawing of the relevant transitions is shown in Fig. 1(b). Details of the experimental setup have been described in Ref. Li et al. (2022). In short, the 88Sr atomic ensemble is produced in a standard two-stage magneto-optical trap (MOT). The dipole-allowed transition at 461 nm and the intercombinaison transition at 689 nm are used for the first and second cooling stages, respectively. At the end of the cooling stages, we have atoms at a temperature of K.
@@
Then, the atoms are transferred into a three-dimensional optical lattice operating at the 813 nm magic wavelength. The optical lattice is generated by a 4.8 W laser (Precilasers system) that is divided into three mutually orthogonal retro-reflected beams that overlap at the atomic cloud. Each lattice beam has a waist of m, and a power of 650(5) mW. The mean trapping frequency is measured to be kHz. To ensure that the lattice beams do not interfere with each other, a frequency difference much larger than the trapping frequency (MHz) is kept among the beams along the three directions. The atoms are confined in the Lamb-Dicke regime ( 0.28). atoms within a cloud of horizontal size m and vertical size m are loaded into the lattice.
@@
A clock laser beam at 698 nm, derived from an external cavity diode laser that is stabilized on an ultra-low expansion cavity (Menlo system), is overlapped with one of the horizontal optical lattice beam using a dichroic mirror [see Fig. 3(a)]. The linewidth of the clock laser system, after frequency stabilization and linewidth narrowing on the cavity, is below 10 Hz. The strontium clock transition is doubly forbidden, and in the case of fermionic isotope, it is weakly allowed thanks to state mixing arising from hyperfine interactions Boyd et al. (2007). As 88Sr has zero nuclear spin, we applied a static magnetic field of 60 G along the polarization axis of the clock laser to activate the clock transition through a state mixing between the and states Taichenachev et al. (2006). The clock laser beam waist is 360 m, making it at least three times larger than the cloud size to minimize inhomogeneous dephasing that might broaden the transition. During the clock-transition interogation, we maintain the power of the clock laser at 5 mW, leading to a Rabi frequency of Hz. In addition, the clock transition experiences a global (quadratic Zeeman and AC-Stark shifts at 698 nm) frequency shift of kHz, which does not affect the magic wavelength position.
@@
A frequency-tunable 140 mW diode laser at 477 nm (Toptica DL pro) is used as a test laser. The beam as a weak ellipticity leading to m, where () is the waist along the -axis. We stabilize the laser frequency on a wavemeter (HighFinesse WS7-30) through its software-based PID controller. To guarantee its ultimate 30 MHz accuracy, the wavemeter is referenced on the 698 nm clock-transition laser Baillard et al. (2007). However, as the accuracy is achieved over a limited spectral range, we transfer it to the blue spectral region using a precise measurement of the 461 nm transition, which we suppose to be close enough to the test laser at 477 nm. We compare the measured frequency of the 461 nm resonance with the value reported in the literature Stellmer and Schreck (2014) with an precision of 5 MHz. Our measured frequency was found to be 146 MHz above the value reported in the literature. This variation, that we arribute to a systematic drift of the wavemeter, is subtracted from each measurement of the 477 nm laser wavelength to guarantee a frequency accuracy that is equal to the accuracy of the wavemeter around the clock transition.
@@
In the presence of the 477 nm test laser, we expect energy shifts of and on the ground and excited states, respectively [see inset of Fig. 3(a)]. If , the atom experiences a non-zero clock transition frequency shift of . To measure this shift in the clock transition frequency, we fit the resulting spectra with a Lorentzian profile to determine the respective line centers. The experimental clock frequency shift is calculated from the frequency difference between the two line centers.
@@ IV Results
We start with a coarse scan of the test laser frequency for a rough determination of corresponding to the magic wavelength condition, namely . Further, we vary the test laser frequency over a range of 10 GHz around , for a finer estimation. The test laser power is set at mW, corresponding to a peak intensity of W/cm2. The value of for each are reported in Fig. 5(a). The experimental results are fitted to a linear function, with the inverse of the error bars acting as weights. This fit is shown as the blue solid line in Fig. 5(a). The intersection point between and the fitting line indicates the frequency of the magic wavelength, which is found to be 628728.2(1) GHz, or equivalently, a wavelength of 476.82362(8) nm in vacuum. The slope is found to be Hz/GHz. The uncertainty mainly comes from the statistical error of the measurement, to which we add the uncertainty of the wavemeter reading.
We compare the experimental results to the theoretical prediction. The frequency of the clock transition experiences a differential AC-Stark shift that reads,
where denotes the differential polarizability between the clock states. is the laser intensity, , and are the vacuum permittivity, and the speed of light in vacuum, respectively. Using the values reported in Sec. II, we estimate the magic wavelength at 476.885(54) nm, which differs by 0.061(54) nm from the experimental value. This small discrepancy might originates from bias in the determination of reduced dipole moments of some transitions, or of the core polarizability.
Additionally, we analyse the variation of the differential AC-Stark shift at the vicinity of the magic wavelength. Within the scanning range of the test laser frequency, we approximate the differential polarizability to be linearly dependent on the frequency detuning . Thus, we have
@@
To test the model validity over an extended parameter range, we measure the frequency shift of the clock line center for various test-laser powers , and frequencies . The results are presented in Fig. 5(b). The plain lines correspond to linear fits whereas the dashed lines are extracted from Eq. (5).
@@ V Conclusion
We have performed an experimental measurement of a blue magic wavelength near 477 nm in 88Sr atoms with a relative frequency uncertainty at the level. The magic wavelength is calculated to be 476.82362(8) nm from a spectroscopic measurement of the clock transition AC-Stark shift. Our experimental value slightly deviates from the theoretical prediction by 0.061(54) nm. Thus, this work could be helpful for a more accurate estimation of the atomic polarizability in the blue spectral region, for example by fine tuning the reduced dipole moment of some transitions or the core contribution.
The 477 nm magic wavelength for the strontium clock transition has the advantage of a shorter wavelength and higher polarizability, compared to the commonly used 813 nm magic wavelength. This makes the 477 nm magic wavelength a good candidate for Bragg pulses to perform matter-wave interferometries involving both the and states, with potential applications for atomic clock interferometry Loriani et al. (2019); Di Pumpo et al. (2021) and quantum test of weak equivalence principle in the optical domain Zych and Brukner (2018); Rosi et al. (2017). However, the photon scattering rate at 477 nm is higher than at 813 nm, making the 477 nm magic wavelength more challenging to use for conventional optical lattice clocks and strongly-correlated gases, two applications that generally require a long coherence time. Nevertheless, cooperative effects in light emission where the timescales involved are generally shorter might be facilitated as the optical lattice could be deeper and accommodate a denser sample.
@@ Acknowledgements.
The authors thank Li Jianing for her contribution to the development of the experimental apparatus. This work was supported by the CQT/MoE (Grant No. R-710-002-016-271), and , the Singapore Ministry of Education Academic Research Fund Tier2 (Grant No. MOE-T2EP50223-0004).References

### v2

@@ Abstract
We report the experimental measurement of a magic wavelength at 476.82362(8) nm for the 88Sr clock transition. The magic wavelength is determined through an AC-Stark shift spectroscopy of atoms in an optical dipole trap. The value is in reasonable agreement with the theoretical prediction, differing by 0.061(54) nm. This magic wavelength, being shorter than the commonly used one at 813 nm, will be important for applications such as Bragg-pulse matter-wave interferometry involving both clock states. This work also paves the way for quantum simulation with a shorter lattice constant.
@@ I Background
Laser-cooled neutral atoms are studied for various applications, such as quantum simulation and computing Bloch et al. (2012); Henriet et al. (2020), precision sensing Geiger et al. (2020), and metrology Ludlow et al. (2015). These applications usually require optical trapping of atoms and coherent control of a transition between two internal states under consideration. Common trapping techniques include large optical dipole traps, optical lattices, and optical tweezers. With sufficiently deep optical lattices and tweezers, it is possible to suppress Doppler broadening and recoil shifts, allowing for high-precision spectroscopy. However, within the trap, spatial inhomogeneity generally reduces the coherence between the states. A common technique to suppress this dephasing mechanism consists in tuning the optical lattice wavelength to the magic wavelength, canceling the differential AC-Stark shift Katori et al. (2003); Takamoto and Katori (2003); Ye et al. (2008).
@@
The magic wavelength value depends on the atomic species and the transition under consideration. Ultra-narrow clock transitions of alkaline-earth atoms have attracted a lot of interest for quantum technological and precision measurement applications. Optical lattice clocks Takamoto et al. (2005); Lemke et al. (2009); Bothwell et al. (2019); Aeppli et al. (2024) that operate on these transitions are among the most precise and accurate clocks and could play an important role in a future redefinition of the second Dimarcq et al. (2024). For the clock transition of strontium atoms at 698 nm, the optical lattice generally operates at a magic wavelength of 813 nm Takamoto and Katori (2003). 813 nm magic-wavelength laser beams have also been used in strontium tweezer array systems Covey et al. (2019); Schine et al. (2022) and quantum gas microscopy Buob et al. (2024), contributing to applications of the clock transition in quantum computation Tsai et al. (2025) and quantum simulation Kolkowitz et al. (2017).
@@
For a given transition, there could be more than one magic wavelength. Finding new magic wavelengths can facilitate some specific usages. For example, the recent measurement of a magic wavelength close to 461 nm for the strontium intercombination line Kestler et al. (2022) led to progress in the state-insensitive trapping of atoms close to a nanofiber Kestler et al. (2023). For the strontium clock transition, we expect three magic wavelengths that are red detuned from the main transition at 461 nm, namely with a positive real part of the polarizability for each clock state. So far, only the magic wavelength at 813 nm has been used, mainly because the photon scattering rate is low, leading to potentially long coherence times. Recently, a shorter magic wavelength at 497 nm has been identified Kestler et al. (2025). Shorter wavelengths can be useful to create an optical lattice with a smaller inter-trap spacing. In addition, magic wavelength Bragg pulses that are used in matter-wave interferometers operating on both clock states Li (2023), can benefit from shorter wavelengths that transfer larger momenta to the atoms.
@@
In this work, we experimentally measure the shortest red-detuned magic wavelength for the strontium clock transition, using the bosonic 88Sr isotope. This magic wavelength, which is at approximately 477 nm, lies between the 474 nm transition and the 483 nm transition [see Fig. 1(a) for the energy level diagram of strontium]. As the closest red-detuned magic wavelength to the 461 nm transition, the atomic polarizability is the largest among the red-detuned magic wavelengths, as shown in Fig. 1(b). The magic wavelength is determined by spectroscopic measurements of the frequency shift of the 698 nm clock transition in the presence of an optical dipole trap. Then, we compare the measured magic wavelength with a theoretical estimation.
@@
This article is organized as follows. In Sec. II, we present the theoretical estimation of the magic wavelength. In Sec. III, we discuss the experimental setup and the procedure for the spectroscopic measurements. In Sec. IV, we report the results on the 477 nm magic-wavelength measurement. In addition, we compare the differential AC-Stark shift measurements with a model that includes the residual inhomogeneous broadening of the clock transition line. In Sec. V, we summarize our studies and provide an outlook for the uses of the 477-nm magic wavelength.
@@ II Polarizabilities estimation
The AC-Stark shift of a state in the electric dipole approximation is given by
@@
Generally, the polarizability can be decomposed into three contributions: the scalar, the vector, and the tensor polarizabilities. However, due to the spherical symmetry of the bosonic strontium clock states, the vector and tensor terms vanish. Thus, the polarizability for state with total angular momentum takes the following polarization-independent form Mitroy et al. (2010),
where is the angular frequency of the bare transition from state to state , is the electric dipole operator, and is the Planck constant. The equation clearly indicates that the transitions with nearby resonant frequencies present the dominant contribution. These transitions are shown in Fig. 1(a) and their respective properties are listed in Table 1. Using Eq. (2) and the values given in Table 1, we compute the polarizabilities of the clock states around 477 nm, as shown in Fig. 2.
@@ III Experiment
The experiment is performed on an ultracold cloud of 88Sr bosonic isotope. A schematic drawing of the relevant transitions is shown in Fig. 1(a). Details of the experimental setup have been described in Ref. Li et al. (2022). In short, the 88Sr atomic ensemble is produced in a two-stage magnetooptical trap (MOT). The dipole-allowed transition at 461 nm and the intercombination transition at 689 nm are used for the first and second cooling stages, respectively. At the end of the cooling stages, we have atoms at a temperature of K.
@@
Then, the atoms are transferred into a three-dimensional optical lattice operating at the 813-nm magic wavelength. The optical lattice is generated by a 4.8 W laser (Precilasers system) that is divided into three mutually orthogonal retro-reflected beams that overlap at the atomic cloud. Each lattice beam has a waist of m, and a power of 650(5) mW. The mean trapping frequency is measured to be kHz. To ensure that the lattice beams do not interfere with each other, a frequency difference of around MHz, much larger than the trapping frequency is kept among the beams along the three directions. The atoms are confined in the Lamb-Dicke regime ( 0.28). atoms within a cloud of horizontal size m and vertical size m are loaded into the lattice.
@@
A clock laser beam at 698 nm, derived from an external cavity diode laser that is stabilized on an ultra-low expansion cavity (Menlo system), is overlapped with one of the horizontal optical lattice beams using a dichroic mirror [see Fig. 3(a)]. The linewidth of the clock laser beam, after frequency stabilization and linewidth narrowing on the cavity, is below 10 Hz. The strontium clock transition is doubly forbidden, and in the case of the fermionic isotope, it is weakly allowed due to state mixing arising from hyperfine interactions Boyd et al. (2007). As 88Sr has zero nuclear spin, we applied a static magnetic field of 60 G along the polarization axis of the clock laser to activate the clock transition through a state mixing between the and states Taichenachev et al. (2006). The clock laser beam waist is 360 m, making it at least three times larger than the cloud size to minimize inhomogeneous dephasing that might broaden the transition. During the clock-transition interrogation, we maintain the power of the clock laser at 5 mW, leading to a Rabi frequency of about Hz. In addition, the clock transition experiences a global (quadratic Zeeman and AC-Stark shifts at 698 nm) frequency shift of about kHz, giving a negligible shift of the magic wavelength position.
@@
A frequency-tunable 140 mW diode laser at 477 nm (Toptica DL pro) is used as a test laser. The beam has a weak ellipticity leading to m, where () is the waist along the -axis. We stabilize the laser frequency on a wavemeter (HighFinesse WS7-30) through its software-based PID controller. To guarantee its ultimate 30 MHz accuracy, the wavemeter is referenced on the 698 nm clock-transition laser Baillard et al. (2007). However, as the accuracy is achieved over a limited spectral range, we transfer it to the blue spectral region, performing an absorption spectroscopy of the 461 nm transition on the ultracold gas. The spectrum is fit to a Lorentzian profile to determine the resonance frequency, which is found to be 146(5) MHz above the value reported in the literature Stellmer and Schreck (2014). This variation, which we attribute to a systematic drift of the wavemeter, is subtracted from each measurement of the 477 nm laser wavelength to ensure a frequency accuracy equal to the accuracy of the wavemeter around the clock transition.
@@
In the presence of the 477 nm test laser, we expect energy shifts of and on the ground and excited states, respectively [see the inset of Fig. 3(a)]. If , the atom experiences a non-zero clock transition frequency shift of . To measure this shift in the clock transition frequency, we fit the resulting spectra with a Lorentzian profile to determine the respective line centers. The experimental clock frequency shift is calculated from the frequency difference between the two line centers.
@@ IV Results
We start with a coarse scan of the test laser frequency for a rough determination of corresponding to the magic wavelength condition, namely . Then, we vary the test laser frequency over a range of 10 GHz around for a finer estimation. The test laser power is set at mW, corresponding to a peak intensity of W/cm2. The values of for each are reported in Fig. 5(a). The experimental results are fit to a linear function, with the inverse of the error bars acting as weights. This fit is shown as the blue solid line in Fig. 5(a). The intersection point between and the fitting line indicates the frequency of the magic wavelength, which is found to be 628728.2(1) GHz, or equivalently, a wavelength of 476.82362(8) nm in vacuum. The slope is found to be Hz/GHz.
Both the linear fit error and the measurement uncertainty of the wavemeter contribute to the overall error of the reported magic wavelength. For the experimental data points in Fig. 5, the error bars are the fit error of the measured spectrum, where statistical error in the fluorescence imaging is the main source of uncertainty. AC-Stark shifts due to the fluctuations and drift of the clock laser power do not contribute significantly to the error, since we perform our measurement in an interleaved manner. Shot-to-shot fluctuation in the bias magnetic field leads to a root mean square frequency fluctuation of 0.5 Hz, smaller than the error bars of . Moreover, we compute the systematic shift of the magic wavelength due to state mixing induced by the bias magnetic field or the 698-nm and 813-nm optical fields. We find a negligible shift below 1 MHz.
We compare the experimental results with the theoretical prediction. The frequency of the clock transition experiences a differential AC-Stark shift given by,
where denotes the differential polarizability between the clock states. is the laser intensity, , and are the vacuum permittivity, and the speed of light in vacuum, respectively. Using the values reported in Sec. II, we estimate the magic wavelength to be 476.885(54) nm, which is in reasonable agreement with the experimental value. The difference of 0.061(54) nm between the two values is almost consistent with zero, lying just outside the error of one standard deviation.
Additionally, we analyze the variation of the differential AC-Stark shift in the vicinity of the magic wavelength. Within the scanning range of the test laser frequency, we approximate the differential polarizability to be linearly dependent on the frequency detuning . Thus, we have
@@
To test the model validity over an extended parameter range, we measure the frequency shift of the clock line center for several values of test-laser peak intensity , and frequencies . The results are presented in Fig. 5(b). The plain lines correspond to linear fits, whereas the dashed lines are extracted from Eq. (5).
@@ V Conclusion
We performed a measurement of a blue magic wavelength near 477 nm in 88Sr atoms with a relative frequency uncertainty at the level. The magic wavelength is found to be 476.82362(8) nm from a spectroscopic measurement of the clock-transition AC-Stark shift. Our experimental value is in reasonable agreement with the theoretical prediction, with a difference of 0.061(54) nm, mainly dominated by the uncertainty in the theoretical calculation. Thus, this work could be helpful for a more accurate estimation of the atomic polarizability in the blue spectral region, for example, by fine-tuning the reduced dipole moment of some transitions or the core contribution.
The 477-nm magic wavelength for the strontium clock transition has the advantage of a shorter wavelength and higher polarizability, compared to the commonly used 813-nm magic wavelength. This makes the 477-nm magic wavelength a good candidate for Bragg pulses to perform matter-wave interferometry involving both the and states, with potential applications for atomic clock interferometry Loriani et al. (2019); Di Pumpo et al. (2021) and quantum test of weak equivalence principle in the optical domain Zych and Brukner (2018); Rosi et al. (2017). However, the photon scattering rate for a given lattice depth is about three orders of magnitude higher at 477 nm than at 813 nm. Hence, the 477-nm magic wavelength is more challenging to use for conventional optical lattice clocks and strongly correlated gases, two applications that generally require a long coherence time. Nevertheless, a denser atomic sample trapped in 477-nm optical tweezers could be useful to study cooperative effects in light emission, since the experimental timescales are generally shorter in these studies.
@@ Acknowledgements.
The authors thank Li Jianing for her contribution to the development of the experimental apparatus. This work was supported by the CQT/MoE (Grant No. R-710-002-016-271) and the Singapore Ministry of Education Academic Research Fund Tier2 (Grant No. MOE-T2EP50223-0004).References
@@
- Mitroy et al. (2010) J Mitroy, M S Safronova, and Charles W Clark, “Theory and applications of atomic and ionic polarizabilities,” Journal of Physics B: Atomic, Molecular and Optical Physics 43, 202001 (2010).